OpenStudy (anonymous):

solve 8-9(2x+4)<-2x+13

OpenStudy (anonymous):

thats the thing i dont get how to

OpenStudy (anonymous):

mby

OpenStudy (anonymous):

ok thnkz

sam (.sam.):

\[\begin{array}{l} 8-9 (2 x+4)<13-2 x \\ \text{expand the terms of the \left hand side} \\ -18 x-28<13-2 x \\ \text{Add }(2 x+28)\text{ \to both sides} \\ -16 x<41 \\\end{array} \\ \\ \text{Divide both sides by negative number flips the sign} \\ \\ x>- \frac{41}{16} \]